# Pushing Cops and Robber on Graphs of Maximum Degree 4

### v1

@@ Abstract.
In a directed graph $\overrightarrow{G}$, the push operation on a vertex $v$ reverses the orientation of all arcs incident on $v$. We consider a variation of classical Cops and Robber on oriented graphs, where in its turn, each cop can either move to an out-neighbor of its current vertex or push some vertex of the graph, whereas, the robber can move to an adjacent vertex in its turn. [Das et al., CALDAM, 2023] introduced this variant and established that if $\overrightarrow{G}$ is an orientation of a subcubic graph, then one cop with push ability has a winning strategy. We extend these results to establish that if $\overrightarrow{G}$ is an orientation of a $3$-degenerate graph, or of a graph with maximum degree $4$, then one cop with push ability has a winning strategy.
@@ 1. Introduction
Cops and Robber is a well-studied pursuit-evasion game, where a set of cops pursue a single robber. We study a variant of Cops and Robber on oriented graphs. Classically, the game in the oriented setting has the following rules. The game starts with the cops placing themselves on the vertices of an oriented graph $\overrightarrow{G}$, and multiple cops may simultaneously occupy the same vertex of the graph. Then the robber chooses a vertex to start. Now the cops and the robber make alternating moves beginning with the cops. In a cop move, each cop can either stay on the same vertex or move to a vertex in its out-neighborhood. In the robber move, the robber does the same. If at some point in the game, one of the cops occupies the same vertex as the robber, we call it the capture. The cops win if they can capture the robber in a finite number of rounds, and if the robber can evade the capture forever, then the robber wins.
The cop number $c(\overrightarrow{G})$ of an oriented graph $\overrightarrow{G}$ is the minimum number of cops needed by the Cop Player to have a winning strategy. We say that an oriented graph $\overrightarrow{G}$ is $k$-copwin if $k$ cops have a winning strategy in $\overrightarrow{G}$. For brevity, we say that $\overrightarrow{G}$ is cop-win if $\overrightarrow{G}$ is $1$-copwin. Most research in oriented (or directed) graphs considers the model defined above. However, there is some research concerning variations of the game in oriented graphs [DGkSS21].
Let $\overrightarrow{uv}$ be an arc of an oriented [^1] An oriented graph is a directed graph without self-loops and 2-cycles. graph $\overrightarrow{G}$. We say that $u$ is an in-neighbor of $v$ and $v$ is an out-neighbor of $u$. Let $N^{-}(u)$ and $N^{+}(u)$ denote the set of in-neighbors and out-neighbors of $u$, respectively. Moreover, let $N^{+}[v]=N^{+}(v)\cup\{v\}$ and $N^{-}[v]=N^{-}(v)\cup\{v\}$. A vertex without any in-neighbor is a source and a vertex without any out-neighbor is a sink. A vertex $v$ is said to be dominating if $N^{+}[v]=V(\overrightarrow{G})$. For a vertex $v$, the push operation on $v$, denoted by $push(v)$, reverses the orientation of each arc incident on $v$. We remark that the push operation is a well-studied modification operation on directed or oriented graphs [FR95, Klo99, KSo98, Mos72, Pre91, Pre86, Pre85, MW00]. In this work, for convenience and for the sake of better readability, we retain the name of an oriented graph even after some vertices have been pushed, allowing a slight abuse of notation. However, there is no scope of confusion to the best of our knowledge.
[DGR+23] initiated the study of Cops and Robber on oriented graphs with respect to the push operation where the players can have the ability to push the vertices of the graph. They defined two kinds of push ability.
@@
Let ${\sf c_{sp}}(\overrightarrow{G})$ be the cop number of $\overrightarrow{G}$ when the cops have the strong push ability and let $\mathsf{c_{wp}}(\overrightarrow{G})$ be the cop number of $\overrightarrow{G}$ when the cops have the weak push ability. It is easy to see that $\mathsf{c_{sp}}(\overrightarrow{G})\leqslant\mathsf{c_{wp}}(\overrightarrow{G})$. Das et al. [DGR+23] also observed that if $\overrightarrow{G}$ is an orientation of a complete multipartite graph, then $c_{sp}(\overrightarrow{G})=1$.
Our Contribution. Das et al. [DGR+23] established that if $\overrightarrow{G}$ is an orientation of a subcubic graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. We begin with observing that their result can be extended to orientations of $3$-degenerate graphs. The main component of the result concerning subcubic graphs was the following result, which shall be useful to us as well.
Proposition 1.1 ([DGR+23]).
@@
Observe that Proposition 1.1 directly implies the following result. We still provide a proof for the sake of completeness.
Theorem 1.2.
@@
We will prove this using contradiction arguments. Consider a minimal graph $\overrightarrow{G}$ such that its underlying graph $G$ is 3-degenerate and ${\sf c_{sp}}(\overrightarrow{G})>1$ (i.e., for every induced subgraph $\overrightarrow{H}$ of $\overrightarrow{G}$, ${\sf c_{sp}}(\overrightarrow{H})=1$). Note that $\overrightarrow{G}$ contains at least two vertices as a single vertex graph is trivially cop-win. Since $G$ is a 3-degenerate graph, there is at least one vertex $v\in V(G)$ such that the degree of $v$ is at most three. Note that in $\overrightarrow{G}$, $|N^{+}(v)|+|N^{-}(v)|\leqslant 3$. Let $\overrightarrow{H}$ be the induced subgraph of $\overrightarrow{G}$ we get after deleting the vertex $v$. By our assumption that $\overrightarrow{G}$ is a minimal graph having strong-push cop number at least two, we have that ${\sf c_{sp}}(\overrightarrow{H})=1$. Then, due to Proposition 1.1, ${\sf c_{sp}}(\overrightarrow{G})=1$, which contradicts our assumption that ${\sf c_{sp}}(\overrightarrow{G})>1$. Therefore, if $\overrightarrow{G}$ is an orientation of a 3-degenerate graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. ∎
Since outerplanar graphs are 2-degenerate [LW70] and 3-dimensional grids are 3-degenerate, we have the following corollary of Theorem 1.2.
Corollary 1.3.
If $\overrightarrow{G}$ is an oriented graph such that its underlying graph $G$ is from one of the following graph classes:
  1. (1)
    outerplanar graphs,
  2. (2)
    subcubic graphs,
  3. (3)
    3-dimensional grids,
  4. (4)
    Apollonian networks,
  5. (5)
    girth 5 planar graphs,
then ${\sf c_{sp}}(\overrightarrow{G}))=1$.
We further extend the result concerning subcubic graphs to establish that if $\overrightarrow{G}$ is an orientation of a graph $G$ with max degree $4$, then ${\sf c_{sp}}(\overrightarrow{G})=1$ in Section 3. We remark that subcubic graphs (and graphs of bounded degree) are well-studied in the case of undirected graphs as well and they are shown to have unbounded cop number [HMGHdlM21] and even are shown to be Meyniel-extremal. To prove our result, we show in Section 3.1 that if $\overrightarrow{G}$ is an orientation of a $4$-regular graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. In particular, we have the following lemma.
@@
Then, using Lemma 1.4 and Theorem 1.2, we obtain the following result.
@@
Related Work. The Cops and Robber game is well studied on both directed and undirected graphs. Hamidoune [Ham87] considered the game on Cayley digraphs. Frieze et al. [FKL12] studied the game on digraphs and gave an upper bound of $\mathcal{O}\left(\frac{n(\log\log n)^{2}}{\log n}\right)$ for cop number in digraphs. Loh and Oh [LO17] considered the game on strongly connected planar digraphs and proved that every $n$-vertex strongly connected planar digraph has cop number $\mathcal{O}(\sqrt{n})$. Moreover, they constructively proved the existence of a strongly connected planar digraph with cop number greater than three, which is in contrast to the case of undirected graphs where the cop number of a planar graph is at most three [AF84]. The computational complexity of determining the cop number of a digraph (and undirected graphs also) is a challenging question in itself. Goldstein and Reingold [GR95] proved that deciding whether $k$ cops can capture a robber is EXPTIME-complete for a variant of Cops and Robber and conjectured that the same holds for classical Cops and Robber as well. Later, Kinnersley [Kin15] proved that conjecture and established that determining the cop number of a graph or digraph is EXPTIME-complete. Kinnersley [Kin18] also showed that $n$-vertex strongly connected cop-win digraphs can have capture time $\Omega(n^{2})$, whereas for undirected cop-win graphs the capture time is at most $n-4$ moves [Gav10].
@@
Hahn and MacGillivray [HM06] gave an algorithmic characterization of the cop-win finite reflexive digraphs and showed that any $k$-cop game can be reduced to $1$-cop game, resulting in an algorithmic characterization for $k$-copwin finite reflexive digraphs. However, these results do not give a structural characterization of such graphs. Darlington et al. [DGGH16] tried to structurally characterize cop-win oriented graphs and gave a conjecture that was later disproved by Khatri et al. [KKKY+19], who also studied the game in oriented outerplanar graphs and line digraphs. Moreover, several variants of the game on directed graphs depending on whether the cop/robber player has the ability to move only along or both along and against the orientations of the arcs are also studied [DGkSS21]. Gahlawat, Myint, and Sen [GMS23] considered these variants under the classical operations like subdivisions and retractions, and also established that all these games are NP-hard.
@@
Recently, the cop number of planar Eulerian digraphs and related families was studied in several articles [dlMHK+21, HM18]. In particular, Hosseini and Mohar [HM18] considered the orientations of integer grid that are vertex-transitive and showed that at most four cops can capture the robber on arbitrary finite quotients of these directed grids. De la Maza et al. [dlMHK+21] considered the straight-ahead orientations of 4-regular quadrangulations of the torus and the Klein bottle and proved that their cop number is bounded by a constant. They also showed that the cop number of every $k$-regularly oriented toroidal grid is at most 13.
@@
Bradshaw et al. [BHT21] proved that the cop number of directed and undirected Cayley graphs on abelian groups has an upper bound of the form of $\mathcal{O}(\sqrt{n})$. Modifying this construction, they obtained families of graphs and digraphs with cop number $\Theta(\sqrt{n})$. The family of digraphs thus obtained has the largest cop number in terms of $n$ of any known digraph construction.
@@ 2. Preliminaries
In this paper, we consider the game on oriented graphs whose underlying graph is simple, finite, and connected. Let $\overrightarrow{G}$ be an oriented graph with $G$ as the underlying undirected graph of $\overrightarrow{G}$. We say that $\overrightarrow{G}$ is an orientation of $G$. We consider the push operation on the vertices of $\overrightarrow{G}$, and hence the orientations of arcs in $\overrightarrow{G}$ might change. So, what we refer to $\overrightarrow{G}$ is the graph with the current orientations. Note that although the orientations of the arcs in $\overrightarrow{G}$ might change, the underlying graph $G$ remains the same. Moreover, it is worth noting that given $\overrightarrow{G}$ and $\overrightarrow{H}$ such that $\overrightarrow{G}$ and $\overrightarrow{H}$ have the same underlying graph, it might be possible that there is no sequence of pushing vertices in $\overrightarrow{G}$ that yields $\overrightarrow{H}$.
@@
Let $v$ be vertex of $\overrightarrow{G}$ and $S$ is a subset of vertices of $\overrightarrow{G}$ (i.e., $S\subseteq V(\overrightarrow{G})$). Then, we say that $v$ is a source in $S$ if $S\subseteq N^{+}[v]$. Moreover, we say that $|N^{+}(v)|$ is the out-degree of $v$, $|N^{-}(v)|$ is the in-degree of $v$, and $|N^{+}(v)|+|N^{-}(v)|$ is the degree of $v$.
The Game. We say that a vertex $v$ is safe if no vertex in $N^{-}[v]$ is occupied by a cop. A vertex $v$ is said to be attacked if there is some cop in $N^{-}(v)$. When we have a single cop, we denote the cop by $\mathcal{C}$. We denote the robber by $\mathcal{R}$ throughout the paper. Let $V^{\prime}\subseteq V(G)$ be a set of vertices. We say that $\mathcal{R}$ is restricted to $V^{\prime}$ if $\mathcal{R}$ cannot move to a vertex $u\in V(G)\setminus V^{\prime}$ without getting captured in the next cop move.
If $\mathcal{R}$ is at a vertex $v$ such that $N^{+}(v)=0$, then we say that $\mathcal{R}$ is trapped at $v$. The following result shall be useful to us.
Proposition 2.1 ([DGR+23]).

### v2

@@ Abstract.
In a directed graph $\overrightarrow{G}$, the push operation on a vertex $v$ reverses the orientation of all arcs incident to $v$. We consider a variation of the classical Cops and Robber on oriented graphs, where in its turn, each cop can either move to an out-neighbor of its current vertex or push some vertex of the graph, whereas, the robber can move to an out-neighbor in its turn. [Das et al., CALDAM, 2023] introduced this variant and established that if $\overrightarrow{G}$ is an orientation of a subcubic graph, then one cop with push ability has a winning strategy. We extend these results to establish that if $\overrightarrow{G}$ is an orientation of a $3$-degenerate graph, or of a graph with maximum degree $4$, then one cop with push ability has a winning strategy. Moreover, we establish that if $\overrightarrow{G}$ can be made to be a directed acyclic graph, then one cop with push ability has a winning strategy.
@@ 1. Introduction
Cops and Robber is a well-studied pursuit-evasion game, where a set of cops pursue a single robber. We study a variant of Cops and Robber on oriented graphs. Classically, the game in the oriented setting has the following rules. The game starts with the cops placing themselves on the vertices of an oriented graph $\overrightarrow{G}$, and multiple cops may simultaneously occupy the same vertex of the graph. Then the robber chooses a vertex to start. Now the cops and the robber make alternating moves, beginning with the cops. In a cop move, each cop can either stay on the same vertex or move to a vertex in its out-neighborhood. In the robber move, the robber does the same. If at some point in the game, one of the cops occupies the same vertex as the robber, we call it the capture. The cops win if they can capture the robber in a finite number of rounds, and if the robber can evade the capture forever, then the robber wins.
The cop number $c(\overrightarrow{G})$ of an oriented graph $\overrightarrow{G}$ is the minimum number of cops needed by the Cop Player to have a winning strategy. We say that $\overrightarrow{G}$ is cop-win if one cop has a winning strategy in $\overrightarrow{G}$. Most research in oriented (or directed) graphs considers the model defined above. However, there has been some research concerning variations of the game in oriented graphs [das2021cops].
Let $\overrightarrow{uv}$ be an arc of an oriented [^1] An oriented graph is a directed graph without self-loops and 2-cycles. graph $\overrightarrow{G}$. We say that $u$ is an in-neighbor of $v$ and $v$ is an out-neighbor of $u$. Let $N^{-}(u)$ and $N^{+}(u)$ denote the set of in-neighbors and out-neighbors of $u$, respectively. Moreover, let $N^{+}[v]=N^{+}(v)\cup\{v\}$ and $N^{-}[v]=N^{-}(v)\cup\{v\}$. A vertex without any in-neighbor is a source, and a vertex without any out-neighbor is a sink. A vertex $v$ is said to be dominating if $N^{+}[v]=V(\overrightarrow{G})$. For a vertex $v$, the push operation on $v$, denoted by $push(v)$, reverses the orientation of each arc incident to $v$. We remark that the push operation is a well-studied modification operation on directed or oriented graphs [fisher-push, klostermeyer1, klostermeyer2, Mosesian, Pretzel1, Pretzel2, Pretzel3, Garyandwood]. In this work, for convenience and for the sake of better readability, we retain the name of an oriented graph even after some vertices have been pushed, allowing a slight abuse of notation. However, there is no scope of confusion to the best of our knowledge.
[das2023cops] initiated the study of Cops and Robber on oriented graphs with respect to the push operation where the players can have the ability to push the vertices of the graph. They defined two kinds of push ability.
@@
In this work we consider oriented graphs whose underlying graphs are finite and connected. In the classical Cops and Robber, we assume connectedness because the cop number of a graph is the sum of the cop numbers of its connected components. Moreover, notice that a disconnected graph cannot be cop-win even if the cop has the strong push ability. Recall that oriented graphs do not admit $2$-cycles. If we allow parallel arcs with opposite orientation, then even for simple graphs such as $4$-cycle with arcs both ways between consecutive vertices, one cop with push ability does not have a winning strategy.
Let ${\sf c_{sp}}(\overrightarrow{G})$ be the cop number of $\overrightarrow{G}$ when the cops have the strong push ability and let $\mathsf{c_{wp}}(\overrightarrow{G})$ be the cop number of $\overrightarrow{G}$ when the cops have the weak push ability. It is easy to see that $\mathsf{c_{sp}}(\overrightarrow{G})\leqslant\mathsf{c_{wp}}(\overrightarrow{G})$. Das et al. [das2023cops] also observed that if $\overrightarrow{G}$ is an orientation of a complete multipartite graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$.
Push operation. Pushing vertices of an oriented graph to obtain an orientation that admits some desirable property is well-studied [klostermeyer1]. In particular, it is NP-complete to decide if, using the push operation, an oriented graph can be made a directed acyclic graph (DAG), a strongly connected digraph, or a semi-connected digraph [klostermeyer1], and if an oriented graph can be made Hamiltonian [klostermeyer2]. Heard and Huang [heard2009kernel] established that it is NP-complete to decide if an oriented graph can be made to contain no directed cycle of odd length using the push operation. Moreover, it is NP-complete to decide if an undirected graph is an underlying graph of an oriented graph whose vertices can be pushed to obtain an oriented clique222an oriented clique is an oriented graph where each pair of non-adjacent vertices is connected by a directed 2-path. [bensmail2017oriented]. Klostermeyer and Šoltés [klostermeyer2000extremal] established that sufficiently large tournaments have high connectivity and can be pushed to obtain an exponential number of Hamiltonian cycles. Pushing operation has been extensively studied in the theory of coloring digraphs and homomorphism of digraphs [sen2017homomorphisms, bensmail2021pushable, bensmail2023pushable].
Special attention has been provided to the problem of deciding if an input oriented graph can be made a DAG using the push operation. Huang, MacGillivray, and Wood [huang2001pushing] characterized, in terms of forbidden subdigraphs, the multipartite tournaments which can be made acyclic using the push operation. Huang, MacGillivray, and Yeo [huang2002pushing] proved that the problem remains NP-complete even when restricted to bipartite digraphs and characterized, in terms of two forbidden subdigraphs, the chordal digraphs and bipartite permutation digraphs which can be made acyclic using the push operation. For the forbidden subdigraph characterization for bipartite permutation graphs, Rizzi [rizzi2006acyclically] provided a polynomial time algorithm.
Observe that if a digraph $\overrightarrow{G}$ can be made a DAG with a single source vertex using the push operation, then ${\sf c_{sp}}(\overrightarrow{G})=1$ as the cop can start at the vertex designated to be the source vertex and push the graph to be a DAG with a single source over the next few rounds. Then, the cop can capture $\mathcal{R}$ in this DAG. We begin with establishing that this observation extends to every oriented graph that can be made a DAG using the push operation. To prove this result, we establish that if we can push a digraph to be a DAG, then we can push it to be a DAG with a single source (in Lemma˜3.2). In particular, we have the following theorem concerning DAGs.
Theorem 1.1.
If $\overrightarrow{G}$ can be made a DAG using the push operation, then ${\sf c_{sp}}(\overrightarrow{G})=1$.
Thus, the class of directed graphs where a cop with strong push ability can win is a superclass of graphs that can be made DAGs using the push operation. We are interested in this class of graphs. Das et al. [das2023cops] established that orientations of subcubic and of interval graphs are in this class of graphs, i.e., if $\overrightarrow{G}$ is an orientation of a subcubic or an interval graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. To establish their result concerning subcubic graphs, Das et al. [das2023cops] proved the following, which shall be useful to us as well.
Proposition 1.2 ([das2023cops]).
@@
The above result argues that removal of vertices with degree at most $3$ does not change the outcome of the game considered similarly to how deletion of corners [^3] A vertex $v$ is a corner if there is a vertex $u$ such that $N[v]\subseteq N[u]$. does not change the outcome of the game in undirected graphs. Observe that Proposition˜1.2 can be used directly to extend the result for subcubic graphs to 3-degenerate graphs. Although it is not mentioned explicitly in [das2023cops], the result should be considered a corollary of their result. We still provide a proof for the sake of completeness.
Corollary 1.3.
@@
Consider a minimal graph $\overrightarrow{G}$ such that its underlying graph $G$ is 3-degenerate and ${\sf c_{sp}}(\overrightarrow{G})>1$ (i.e., for every proper induced subgraph $\overrightarrow{H}$ of $\overrightarrow{G}$, ${\sf c_{sp}}(\overrightarrow{H})=1$). Note that $\overrightarrow{G}$ contains at least two vertices as a single vertex graph is trivially cop-win. Since $G$ is a 3-degenerate graph, there is at least one vertex $v\in V(\overrightarrow{G})$ such that $|N^{+}(v)|+|N^{-}(v)|\leqslant 3$. Let $\overrightarrow{H}$ be the induced subgraph of $\overrightarrow{G}$ we get after deleting the vertex $v$. Since $\overrightarrow{G}$ is a minimal graph with ${\sf c_{sp}}(\overrightarrow{G})>1$, we have that ${\sf c_{sp}}(\overrightarrow{H})=1$. Then, due to Proposition˜1.2, ${\sf c_{sp}}(\overrightarrow{G})=1$, which contradicts our assumption that ${\sf c_{sp}}(\overrightarrow{G})>1$. Therefore, if $\overrightarrow{G}$ is an orientation of a 3-degenerate graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. ∎
One can obtain using Euler’s formula that triangle-free planar graphs are 3-degenerate. Moreover, outerplanar graphs are 2-degenerate [outDegenerate] and 3-dimensional grids are 3-degenerate. Thus, Corollary˜1.3 immediately adds these graph classes to the list of graph classes whose orientations are cop-win if the cop has the strong push ability.
We further extend the result concerning subcubic graphs to establish that if $\overrightarrow{G}$ is an orientation of a graph $G$ with maximum degree $4$, then ${\sf c_{sp}}(\overrightarrow{G})=1$ in Section˜4. We remark that subcubic graphs (and graphs of bounded degree) are well-studied in the case of undirected graphs as well and they are shown to have unbounded cop number [hosseini2021meyniel] and even are shown to be Meyniel-extremal [^4] Meyniel’s conjecture [frankl1987cops] asserts that the cop number of any undirected graph is $\mathcal{O}(\sqrt{n})$. A graph class is said to be Meyniel-extremal if it contains some graphs with cop number $\Omega(\sqrt{n})$.. To prove our result, we establish (in Section˜4) that if $\overrightarrow{G}$ is an orientation of a $4$-regular graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. In particular, we have the following lemma.
@@
Then, using Lemma˜1.4, Proposition˜1.2, and Corollary˜1.3, we obtain the following result.
@@
Related Work. The Cops and Robber game is well studied on both directed and undirected graphs. Hamidoune [hamidoune] considered the game on Cayley digraphs. Frieze et al. [lohjgt] studied the game on digraphs and gave an upper bound of $\mathcal{O}\left(\frac{n(\log\log n)^{2}}{\log n}\right)$ for cop number in digraphs. Loh and Oh [loh] considered the game on strongly connected planar digraphs and proved that every $n$-vertex strongly connected planar digraph has cop number $\mathcal{O}(\sqrt{n})$. Moreover, they constructively proved the existence of a strongly connected planar digraph with cop number greater than three, which is in contrast to the case of undirected graphs where the cop number of a planar graph is at most three [fromme1984game]. The computational complexity of determining the cop number of a digraph (and undirected graphs also) is a challenging question in itself. Goldstein and Reingold [exptime] proved that deciding whether $k$ cops can capture a robber is EXPTIME-complete for a variant of Cops and Robber and conjectured that the same holds for the classical Cops and Robber as well. Later, Kinnersley [kinnersley2015cops] proved that conjecture and established that determining the cop number of a graph or digraph is EXPTIME-complete. Kinnersley [kinnersley2018bounds] also showed that $n$-vertex strongly connected cop-win digraphs can have capture time $\Omega(n^{2})$, whereas for undirected cop-win graphs the capture time is at most n4n-4 moves [gavenvciak2010cop].
@@
Hahn and MacGillivray [hahn] gave an algorithmic characterization of the cop-win finite reflexive digraphs and showed that any $k$-cop game can be reduced to $1$-cop game, resulting in an algorithmic characterization for $k$-copwin finite reflexive digraphs. However, these results do not give a structural characterization of such graphs. Darlington et al. [darlington2016cops] tried to structurally characterize cop-win oriented graphs and gave a conjecture that was later disproved by Khatri et al. [khatri2018study], who also studied the game in oriented outerplanar graphs and line digraphs. Moreover, several variants of the game on directed graphs depending on whether the cop/robber player has the ability to move only along or both along and against the orientations of the arcs are also studied [das2021cops]. Gahlawat, Myint, and Sen [gahlawat2023cops] considered these variants under the classical operations like subdivisions and retractions, and also established that all these games are NP-hard even for $2$-degenerate bipartite graphs with arbitrary high girth.
@@
Recently, the cop number of planar Eulerian digraphs and related families was studied in several articles [mohar2, mohar]. In particular, Hosseini and Mohar [mohar] considered the orientations of integer grid that are vertex-transitive and showed that at most four cops can capture the robber on arbitrary finite quotients of these directed grids. De la Maza et al. [mohar2] considered the straight-ahead orientations of 4-regular quadrangulations of the torus and the Klein bottle and proved that their cop number is bounded by a constant. They also showed that the cop number of every $k$-regularly oriented toroidal grid is at most 13.
@@
Bradshaw et al. [bradshaw2021cops] proved that the cop number of directed and undirected Cayley graphs on abelian groups is $\mathcal{O}(\sqrt{n})$, and constructed families of graphs and digraphs which are Cayley graphs of abelian groups with cop number $\Theta(\sqrt{n})$ for every $n$. The family of digraphs thus obtained has the largest cop number in terms of $n$ of any known digraph construction.
@@ 2. Preliminaries
In this paper, we consider the game on oriented graphs whose underlying graph is simple, finite, and connected. Let $\overrightarrow{G}$ be an oriented graph with $G$ as the underlying undirected graph of $\overrightarrow{G}$. We say that $\overrightarrow{G}$ is an orientation of $G$. We consider the push operation on the vertices of $\overrightarrow{G}$, and hence the orientations of arcs in $\overrightarrow{G}$ might change. We generally use $\overrightarrow{G}$ to refer to the graph G with its orientation at some specific point in the game. Note that although the orientations of the arcs in $\overrightarrow{G}$ might change, the underlying graph $G$ remains the same. Moreover, it is worth noting that given $\overrightarrow{G}$ and $\overrightarrow{H}$ such that $\overrightarrow{G}$ and $\overrightarrow{H}$ have the same underlying graph, it might be possible that there is no sequence of pushing vertices in $\overrightarrow{G}$ that yields $\overrightarrow{H}$.
@@
Let $v$ be a vertex of $\overrightarrow{G}$ and $S$ be a subset of vertices of $\overrightarrow{G}$ (i.e., $S\subseteq V(\overrightarrow{G})$). Then, we say that $v$ is a source in $S$ if $S\subseteq N^{+}[v]$. Moreover, we say that $|N^{+}(v)|$ is the out-degree of $v$, $|N^{-}(v)|$ is the in-degree of $v$, and $|N^{+}(v)|+|N^{-}(v)|$ is the degree of $v$. A vertex $w$ is said to be reachable from $u$ if either $w=u$ or there is a directed path from $u$ to $w$.
When we have a single cop, we denote the cop by $\mathcal{C}$. We denote the robber by $\mathcal{R}$ throughout the paper. If $\mathcal{R}$ is at a vertex $v$ such that $N^{+}(v)=0$, then we say that $\mathcal{R}$ is trapped at $v$. The following result shall be useful to us.
Proposition 2.1 ([das2023cops]).
@@
3. Graphs Pushable to be DAGs
In this section, we establish that if $\overrightarrow{G}$ can be made a DAG using the push operations, then ${\sf c_{sp}}(\overrightarrow{G})=1$. To this end, first we prove the following lemma.
Lemma 3.1.
Let $\overrightarrow{G}$ be a DAG and $u$ be a source vertex of $\overrightarrow{G}$. Moreover, let $X$ and $Y\neq\emptyset$ be the sets of vertices that are reachable from $u$ and not reachable from $u$, respectively. Then, we can push some vertices of $Y$ such that:
  1. (1)
    $\overrightarrow{G}$ is still a DAG and $u$ is still a source vertex in $\overrightarrow{G}$,
  2. (2)
    each vertex in $X$ is still reachable from $u$,
  3. (3)
    and at least one vertex in $Y$ is reachable from $u$.
Proof.
For any $x\in X$ and $y\in Y$, clearly $\overrightarrow{xy}\notin E(\overrightarrow{G})$ as otherwise $y$ is also reachable from $u$ due to transitivity of reachability. Let $Y^{\prime}$ be the set of vertices in $Y$ that has at least one out-neighbor in $X$. Observe that $Y^{\prime}\neq\emptyset$ since the underlying graph of $\overrightarrow{G}$ is connected. Now, push each vertex in $Y$. Notice that this operation reverses the orientation of each arc with one endpoint in $X$ and one endpoint in $Y^{\prime}$, ensuring that each vertex in $Y^{\prime}$ is now reachable from $u$. Every other arc has the same orientation still. Thus, each vertex in $X$ is reachable from $u$ and each vertex in $Y^{\prime}$ is reachable from $u$ (and $|Y^{\prime}|>0$). Next, observe that $u$ is still a source vertex since no arc incident on $u$ changed its orientation. Finally, observe that $\overrightarrow{G}$ is still a DAG after pushing vertices of $Y$. Indeed, there can be no directed cycle that contains at least one vertex from $X$ and at least one vertex from $Y$ as all arcs between vertices of $X$ and $Y$ are oriented from $X$ to $Y$, and if there is a cycle in $X$ or in $Y$, it would imply that $\overrightarrow{G}$ contained a cycle before pushing $Y$, contradicting the fact that $\overrightarrow{G}$ was a DAG. This completes our proof. ∎
Using Lemma˜3.1, we establish that if $\overrightarrow{G}$ is a DAG, then $\overrightarrow{G}$ can be pushed to be a DAG with a single source in the following lemma.
Lemma 3.2.
If $\overrightarrow{G}$ is a DAG, then $\overrightarrow{G}$ can be made a DAG with a single source using the push operations.
Proof.
Since $\overrightarrow{G}$ is a DAG, it contains at least one source vertex, say $u$. If every vertex in $\overrightarrow{G}$ is reachable from $u$, then we have nothing to prove as $\overrightarrow{G}$ is a DAG with a single source. Otherwise, let $Z$ be the set of vertices that are not reachable from $u$. Then, we apply Lemma˜3.1 iteratively to ensure that after each application of Lemma˜3.1, at least one vertex of $Z$ that was not reachable from $u$ before this application of Lemma˜3.1 is now reachable from $u$ while ensuring that (i) $\overrightarrow{G}$ remains a DAG, (ii) $u$ remains a source vertex, and (iii) every vertex that was reachable in a previous round is still reachable. Since $Z\subseteq V(\overrightarrow{G})$ is finite and at least one vertex from $Z$ that was not reachable before becomes reachable in each application of Lemma˜3.1, after finitely many applications of Lemma˜3.1, $\overrightarrow{G}$ will be a DAG with a single source $u$. ∎
Finally, we have the following theorem. See 1.1
Proof.
First, push the vertices of $\overrightarrow{G}$ such that now $\overrightarrow{G}$ is oriented to be a DAG. Second, apply Lemma˜3.2 to orient $\overrightarrow{G}$ as a DAG with a single source using the push operations. Now, let $u$ be the unique source of $\overrightarrow{G}$. Since DAGs with a unique source are cop-win when $\mathcal{R}$ has to follow the orientations of the arcs [das2021cops], the cop has a winning strategy in $\overrightarrow{G}$ by starting the game at $u$, pushing the vertices of $\overrightarrow{G}$ to make it a DAG with a unique source $u$, and then capturing the robber using its winning strategy for DAGs with a single source. This completes our proof. ∎